Find the solution for the following equation:
-3/5=x-3/10

Answer by ewatrrr(24785)


Hi
Find the solution for the following equation:
-3/5=x-3/10
-3/5 + 3/10 = x
-6/10 + 3/10 = x
-3/10 = x